# thanks: This paper has been submitted in conjunction with its companion Algeri et al. [1]

### v1

@@ III.2 Distribution-free testing of the angular power spectum
$\hat{\bm{\epsilon}}_{{f_{b}},s}=\tilde{\bm{A}}_{{f_{b}},s}-\tilde{\bm{A}}_{{f_{b}},s}(\hat{\bm{\theta}}).$ (17)
In principle, one could test the hypotheses in Eq. (LABEL:eqn:test) using classical goodness-of-fit statistics constructed as the sum of some function of the sphered residuals – with Pearson’s $\chi^{2}$ being possibly the most prominent example among those – and derive their distribution under the hypothesis $H_{0}$ using the parametric bootstrap [Cf. 41] based on the generalized $\chi^{2}$ distribution derived in Appendix A. Such statistics, however, are known to exhibit no statistical power [^2] In statistics, the concept of ‘power’ refers to the sensitivity of a test in detecting departures from the model hypothesized under the null hypothesis. toward infinitely many possible ‘weak’ deviations from the postulated model [e.g., 42, 43, 44]. That is, there are infinite possible ways in which the hypothesized model could be wrong, but such statistics would not be able to identify them. Test statistics based on the process of partial sums of the residuals [e.g., 45, 46], however, can restore statistical power [e.g., 42, 43, 44] and are thus recommended.
@@ III.2.1 Model testing via partial sums of the residuals
$\bm{\epsilon}=\left[\bm{\epsilon}_{f_{1},1}^{T}(\bm{\theta}_{0}),\dots,\bm{\epsilon}_{f_{B},1}^{T}(\bm{\theta}_{0}),\dots,\bm{\epsilon}_{f_{1},S}^{T}(\bm{\theta}_{0}),\dots,\bm{\epsilon}_{f_{B},S}^{T}(\bm{\theta}_{0})\right]^{T},$
@@
$\hat{\bm{\epsilon}}=\left[\hat{\bm{\epsilon}}_{f_{1},1}^{T},\dots,\hat{\bm{\epsilon}}_{f_{B},1}^{T},\dots,\hat{\bm{\epsilon}}_{f_{1},S}^{T},\dots,\hat{\bm{\epsilon}}_{f_{B},S}^{T}\right]^{T}.$
The process of partial sums of the sphered residuals is
@@
$\begin{split}\hat{\bm{\epsilon}}_{{f_{b}},s}=&\bm{\epsilon}_{{f_{b}},s}(\bm{\theta}_{0})\\ &-\frac{1}{N}\dot{\bm{A}}_{{f_{b}},s}(\bm{\theta}_{0})\Bigl{[}\bm{R}_{p}^{-1}\sum_{{f_{b}},s}\dot{\bm{A}}_{{f_{b}},s}(\bm{\theta}_{0})^{T}\bm{\epsilon}_{{f_{b}},s}(\bm{\theta}_{0})\Bigl{]}+o_{p}(1).\\ \end{split}$ (21)
@@
$\dot{\bm{A}}(\bm{\theta}_{0})=\begin{bmatrix}\dot{\bm{A}}^{T}_{f_{1},1}(\bm{\theta}_{0}),\cdots,\dot{\bm{A}}^{T}_{f_{B},1}(\bm{\theta}_{0}),\cdots,\dot{\bm{A}}^{T}_{f_{B},S}(\bm{\theta}_{0})\end{bmatrix}^{T},$
@@
where $\bm{\mu}_{\bm{\theta}_{j}}$ corresponds to the $jth$ column of the matrix
$\frac{1}{\sqrt{N}}\dot{\bm{A}}(\bm{\theta}_{0})\bm{R}_{p}^{-1/2}.$
@@
The process in Eq. (19) can be shown to be asymptotically distributed as a Brownian motion if $H_{0}$ in Eq. (LABEL:eqn:test) is true. This result follows from the fact that such a process consists of partial sums of weakly dependent [^3] Intuitively, a sequence of random variables is said to be ‘weakly dependent’ if the dependence among the elements of the sequence reduces and eventually ceases to exist for elements that are sufficiently far away from one another within the sequence. quantities [Cf. 48, 49] with mean zero and finite variance. The condition of weak dependence is satisfied since the subvectors $\bm{\epsilon}_{{f_{b}},s}(\bm{\theta}_{0})$ of $\bm{\epsilon}$ are all independent from one another. It follows that since $w_{k,N}$ is a projection of the process in Eq. (19), its limiting distribution under $H_{0}$ is that of a projected Brownian motion. Its mean and covariance are
@@
Consider the sphered errors generated either from a standard multivariate Gaussian distribution or independently from Laplace distribution with mean 00 and variance $1$. Let the covariance matrices $\bm{\Sigma}_{{f_{b}},s}$ be generated from a Wishart distribution with 10 degrees of freedom, and these matrices are assumed to be known. Denote the unsphered errors as
@@
for all $s=1,\dots,S$. We consider four datasets generated from the four possible additive combinations of models and error terms in Eqs. (26)-(27) with $\bm{\theta}=(4,$3,0.4)3,0.4)3 , 0.4 ). For each of these datasets, we perform $S=15$ repeated simulations, resulting in an effective sample size of $120,$ and a total of $N=960$. The datasets are generated $10^{4}$ times; each time, we calculate a single realization of the process $w_{k,N}$ in Eq. (23), which yields one observation of the statistics $K_{N}$ and $C_{N}$ from Eq. (25).
@@ III.2.2 Construction of distribution-free test statistics
There are many possible ways to choose the orthonormal set $\{\bm{r}_{j}\}_{j=1}^{p}$. For example, as proposed by Khmaladze [47], one could consider
$\bm{r}_{1}=(1/\sqrt{N},\dots,1/\sqrt{N})^{T},$
letting $\bm{r}_{2}$ be the vector of elements
$\bm{r}_{2,n}=\sqrt{\frac{12N}{N^{2}-1}}\left(\frac{n}{N}-\frac{N+1}{2N}\right),\ \ n=1,\ldots,N;$ (30)
the remaining vectors $\bm{r}_{3},\ldots,\bm{r}_{p}$ can be obtained by the Gram–Schmidt process applied to the powers of $\bm{r}_{2}$. Alternatively, one could construct the orthonormal set $\{\bm{r}_{j}\}_{j=1}^{p}$ by considering a set of $p$ orthonormal functions with a given inner product. For example, the Hermite polynomials $\{H_{j}(x)\}_{j=1}^{p}$ satisfy the condition:
[…]
We can then specify the elements of the vectors $\bm{r}_{j}$ as
$r_{j,n}=\frac{1}{\sqrt{c_{j}}}H_{j}(x_{n})\sqrt{e^{-x_{n}^{2}}},$
where $c_{j}=\sum_{n}H^{2}_{j}(x_{n})e^{-x_{n}^{2}}$ are normalizing constants and $x_{n}=-\zeta+\frac{2\zeta(n-1)}{N-1},$ $n=1,\dots,N$, represent $N$ equally spaced points over the interval $[-\zeta,\zeta]$ for some appropriately chosen constant $\zeta>0$.
@@
$U_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}$ $\bm{\mu}_{\bm{\theta}_{1}}=\bm{r}_{1}$
[…]
@@
$\bm{U}_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}\bm{U}_{\bm{\mu}_{\bm{\theta}_{2}},\bm{r}_{2}}\bm{\mu}_{\bm{\theta}_{2}}=\bm{U}_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}\bm{r}_{2}\neq\bm{r}_{2}.$
@@
$\langle\widetilde{\bm{r}}_{2},\bm{\mu}_{\bm{\theta}_{1}}\rangle$ $=\langle\bm{r}_{2},U_{\bm{\mu}_{\bm{\theta}_{1}},\widetilde{\bm{r}}_{1}}\bm{\mu}_{\bm{\theta}_{1}}\rangle=\langle\bm{r}_{2},\bm{r}_{1}\rangle=0,$
@@ III.3.1 Distribution-free Property
To demonstrate the validity of the distribution-free property of the method outlined in Section III.2.2, let us revisit the example introduced in Section III.2. Under the same setup, we simulate the null distributions of the statistics $\widetilde{K}_{N}$ and $\widetilde{C}_{N}$ in Eq. (29) with $v_{k,N}$ constructed as in Eq. (39) for each of the four combinations of models and error structures given in Eqs. (26)-(27). For the sake of comparison, we also simulate the null distribution of $\widetilde{K}_{N}$ and $\widetilde{C}_{N}$ using the Monte Carlo procedure described at the end of Section III.2.2 and for which the errors are generated from a standard Gaussian. Figure 2 shows that regardless of the error structure or model considered, the graphs of the simulated null distributions overlap for both the Kolmogorov-Smirnov and the Cramér-Von Mises statistics. This confirms that the proposed tests are, for sufficiently large $N$, distribution-free, and the proposed Monte Carlo scheme accurately approximates the limiting null distributions of the test statistics based on $v_{k,N}$. The CPU time required by the latter amounts to 3.74 seconds.
@@ III.3.2 Statistical properties
To ensure that the probability of a false rejection – known in statistics as the probability of Type I error – does not exceed the pre-determined significance level, we generate $10^{4}$ datasets from the additive combination of the mean function $\bm{A}^{M1}_{{f_{b}},\ell}(\bm{\theta})$ and the Gaussian error $\bm{u}^{(1)}_{{f_{b}},s}$ in Eqs. (26)-(27). We test the validity of $\bm{A}^{M1}_{{f_{b}},\ell}(\bm{\theta})$ by means of the statistics $\widetilde{K}_{N}$ and $\widetilde{C}_{N}$. Table 1 shows the simulated probabilities of rejecting $\bm{A}^{M1}_{{f_{b}},\ell}(\bm{\theta})$ when choosing different significance levels. The results indicate that the simulated Type I error probabilities closely match the nominal levels.
@@
A Monte Carlo simulation involving $10^{4}$ replicates is performed for each of the models in Eq. (42) and combined with the Gaussian error structure in Eq. (27). The simulated power of the Kolmogorov–Smirnov and Cramér–von Mises test statistics are reported in Tables 2 and 3, respectively, for different significance levels $\alpha$. For both test statistics, the simulated statistical power is close to one when the significance level is 0.05 and is reasonably high even at smaller significance levels.
@@ IV Testing spectral models for the SGWB
We now apply the methodology described in Sec. III to the case of a realistic astrophysical SGWB model and the data obtained by the Advanced LIGO and Advanced Virgo GW detectors in their third observing run (O3), as described in Sec. II.
@@
[…]
@@
$\mathcal{B}(f,z;\bm{\theta})=A_{\text{max}}\,f^{-1/3}\,e^{-(z-z_{c})^{2}/2\sigma_{z}^{2}}\,,$ (47)
with the free parameters $\theta=(A_{\text{max}},z_{c},\sigma_{z})$.
@@
$\mathcal{K}(f)=\frac{2\pi^{2}}{3H_{0}^{2}}f_{ref}^{3}\left(\frac{f}{f_{ref}}\right)^{\alpha}$ (51)
Here, $f_{ref}$ is the reference frequency of 25Hz used in O3 [7], and $f$ again takes on values of the midpoints of each frequency band $\bar{{f_{b}}}$. We also apply the square of this scaling factor to the inverted Fisher matrix in order to obtain a properly scaled covariance matrix for the $A_{\ell}$ (Eq. (7)). For notational simplicity, for the remainder of this section, when we write ${a}_{{f_{b}},\ell m}^{M}$, we assume it to be the version scaled by $\tau_{{f_{b}},\ell}$, and when we write $\hat{a}_{{f_{b}},s,\ell m}$ or $(\Gamma^{-1}_{{f_{b}},s})_{\ell m,\ell^{\prime}m^{\prime}}$, we assume them to be the versions scaled by $\mathcal{K}(f)$ and $\mathcal{K}(f)^{2}$ respectively.
With the model map elements ${{a}}_{{f_{b}},\ell m}^{M}$ and data map elements $\hat{{a}}_{{f_{b}},s,\ell m}$ properly scaled, we may find their sum ${\hat{a}}_{{f_{b}},s,\ell m}^{inj}$ to complete the signal injection. The corresponding angular power spectrum ${\hat{A}}_{{f_{b}},s,\ell}^{inj}$ is calculated according to Eq. (6),
$\hat{A}^{inj}_{{f_{b}},s,\ell}=\frac{1}{1+2\ell}\sum_{m=-\ell}^{\ell}\left[|\hat{a}^{inj}_{{f_{b}},s,\ell m}|^{2}-(\Gamma^{-1}_{{f_{b}},s})_{\ell m,\ell m}\right],$ (52)
and the covariance matrix is computed according to Eq. (7), with ${a}_{{f_{b}},\ell m}$ replaced by $\hat{a}^{inj}_{{f_{b}},\ell m}$, which is defined as the average of $\hat{a}^{inj}_{{f_{b}},s,\ell m}$ over the data segments $s$.
The necessary inputs for the procedure described in Sec. III are the data ${\hat{A}}_{{f_{b}},s,\ell}^{inj}$, their corresponding covariance matrices, and model values ${A}_{\ell}(\bar{{f_{b}}};\bm{\theta})$. In this study, we test our procedure on two variations of these quantities.
We compare the model described by Eqs. (45)-(47) to data injected using the same model, applying the covariance matrix as described above. When generating the injected signal, the parameter values are $1.2\times 10^{-28}$, $0.6$, and $0.7$ for $A_{max}$, $z_{c}$, and $\sigma_{z}$ respectively. The parameter estimates obtained via Eq (11) are $1.19\times 10^{-28}$, $0.590$, and $0.694$. Next, we use the same covariance matrix and again use a postulated model as given in Eq. (47), but in this case we use data which has been injected with a different variation of the model. In particular, we replace the astrophysical kernel with
@@
where $\bm{\theta}=(\beta_{0},\beta_{1})$. For this alternative model, we set $\beta_{0}=1.2\times 10^{-28}$ and $\beta_{1}=6.0\times 10^{-29}$ when generating the injected data. The parameter estimates for the postulated model are then $2.84\times 10^{-28}$, $3.58$, and $2.72$ for $A_{max}$, $z_{c}$, and $\sigma_{z}$, respectively.

### v3

@@ III.2 Distribution-free testing of the angular power spectum
[…]
In principle, one could test the hypotheses in Eq. (LABEL:eqn:test) using classical goodness-of-fit statistics constructed as the sum of some function of the sphered residuals – with Pearson’s $\chi^{2}$ being possibly the most prominent example among those – and derive their distribution under the hypothesis $H_{0}$ using the parametric bootstrap [Cf. 47] based on the generalized $\chi^{2}$ distribution derived in Appendix A. Such statistics, however, are known to exhibit no statistical power [^6] In statistics, the concept of ‘power’ refers to the sensitivity of a test in detecting departures from the model hypothesized under the null hypothesis. toward infinitely many possible ‘weak’ deviations from the postulated model [e.g., 48, 49, 50]. That is, there are infinite possible ways in which the hypothesized model could be wrong, but such statistics would not be able to identify them. Test statistics based on the process of partial sums of the residuals [e.g., 51, 52], however, can restore statistical power [e.g., 48, 49, 50] and are thus recommended.
@@ III.2.1 Model testing via partial sums of the residuals
$\bm{\epsilon}=\left[\bm{\epsilon}_{f_{1},1}^{T}(\bm{\theta}_{0}),\dots,\bm{\epsilon}_{f_{B},1}^{T}(\bm{\theta}_{0}),\dots,\bm{\epsilon}_{f_{1},S}^{T}(\bm{\theta}_{0}),\dots,\bm{\epsilon}_{f_{B},S}^{T}(\bm{\theta}_{0})\right]^{T},$ (29)
@@
$\hat{\bm{\epsilon}}=\left[\hat{\bm{\epsilon}}_{f_{1},1}^{T},\dots,\hat{\bm{\epsilon}}_{f_{B},1}^{T},\dots,\hat{\bm{\epsilon}}_{f_{1},S}^{T},\dots,\hat{\bm{\epsilon}}_{f_{B},S}^{T}\right]^{T}.$ (30)
The process [^7] The term “process” here denotes a stochastic process, where $w_{k,N}$ is a sequence of discrete random variables indexed by $k$ and of length $N$. of partial sums of the sphered residuals is
@@
[…]
@@
$\dot{\bm{A}}(\bm{\theta}_{0})=\begin{bmatrix}\dot{\bm{A}}^{T}_{f_{1},1}(\bm{\theta}_{0}),\cdots,\dot{\bm{A}}^{T}_{f_{B},1}(\bm{\theta}_{0}),\cdots,\dot{\bm{A}}^{T}_{f_{B},S}(\bm{\theta}_{0})\end{bmatrix}^{T},$ (35)
@@
where $\bm{\mu}_{\bm{\theta}_{j}}$ corresponds to the jthjth column of the matrix
$\frac{1}{\sqrt{N}}\dot{\bm{A}}(\bm{\theta}_{0})\bm{R}_{p}^{-1/2}.$ (37)
@@
The process in Eq. (32) can be shown to be asymptotically distributed as a Brownian motion if $H_{0}$ in Eq. (LABEL:eqn:test) is true. This result follows from the fact that such a process consists of partial sums of weakly dependent [^8] Intuitively, a sequence of random variables is said to be ‘weakly dependent’ if the dependence among the elements of the sequence reduces and eventually ceases to exist for elements that are sufficiently far away from one another within the sequence. quantities [Cf. 54, 55] with mean zero and finite variance. The condition of weak dependence is satisfied since the subvectors $\bm{\epsilon}_{{f_{b}},s}(\bm{\theta}_{0})$ of $\bm{\epsilon}$ are all independent from one another. It follows that since $w_{k,N}$ is a projection of the process in Eq. (32), its limiting distribution under $H_{0}$ is that of a projected Brownian motion. Its mean and covariance are
@@
Consider the sphered errors generated either from a standard multivariate Gaussian distribution or independently from Laplace distribution with zero mean and unit variance. Let the covariance matrices $\bm{\Sigma}_{{f_{b}},s}$ be generated from a Wishart distribution with 10 degrees of freedom, and these matrices are assumed to be known. Denote the unsphered errors as
@@
for all $s=1,\dots,S$. We consider four datasets generated from the four possible additive combinations of models and error terms in Eqs. (41)-(42) with $\bm{\theta}=(4,$ -$3,0.4)$. For each of these datasets, we perform $S=15$ repeated simulations, resulting in an effective sample size of $120,$ and a total of $N=960$. The datasets are generated $10^{4}$ times; each time, we calculate a single realization of the process $w_{k,N}$ in Eq. (38), which yields one observation of the statistics $K_{N}$ and $C_{N}$ from Eq. (40).
@@ III.2.2 Construction of distribution-free test statistics
$U_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}$ $\bm{\mu}_{\bm{\theta}_{1}}=\bm{r}_{1}$ (49)
$\text{ and }\ \ \bm{\mu}_{\bm{\theta}_{1}}^{T}\bm{\epsilon}=\langle U_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}$ $\bm{\mu}_{\bm{\theta}},U_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}\bm{\epsilon}\rangle=\bm{r}_{1}^{T}\bm{e}.$
@@
$\bm{U}_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}\bm{U}_{\bm{\mu}_{\bm{\theta}_{2}},\bm{r}_{2}}\bm{\mu}_{\bm{\theta}_{2}}=\bm{U}_{\bm{\mu}_{\bm{\theta}_{1}},\bm{r}_{1}}\bm{r}_{2}\neq\bm{r}_{2}.$ (52)
@@
$\langle\widetilde{\bm{r}}_{2},\bm{\mu}_{\bm{\theta}_{1}}\rangle$ $=\langle\bm{r}_{2},U_{\bm{\mu}_{\bm{\theta}_{1}},\widetilde{\bm{r}}_{1}}\bm{\mu}_{\bm{\theta}_{1}}\rangle=\langle\bm{r}_{2},\bm{r}_{1}\rangle=0,$ (56)
@@ III.3.1 Distribution-free Property
To demonstrate the validity of the distribution-free property of the method outlined in Section III.2.2, let us revisit the example introduced in Section III.2. Under the same setup, we simulate the null distributions of the statistics $\widetilde{K}_{N}$ and $\widetilde{C}_{N}$ in Eq. (44) with $v_{k,N}$ constructed as in Eq. (59) for each of the four combinations of models and error structures given in Eqs. (41)-(42). We adopt the orthonormal basis $\{\bm{r}_{j}\}_{j=1}^{p}$ constructed following Khmaladze [53], where $\bm{r}_{1}=(1/\sqrt{N},\dots,1/\sqrt{N})^{T}$ and $\bm{r}_{2}$ are the vector of elements
$\bm{r}_{2,n}=\sqrt{\frac{12N}{N^{2}-1}}\left(\frac{n}{N}-\frac{N+1}{2N}\right),\ \ n=1,\ldots,N.$ (62)
The remaining vectors $\bm{r}_{3},\ldots,\bm{r}_{p}$ are obtained by applying the Gram–Schmidt procedure to successive powers of $\bm{r}_{2}$. This orthonormal basis are also adopted for the following simulation studies and the data analysis.
For the sake of comparison, we also simulate the null distribution of $\widetilde{K}_{N}$ and $\widetilde{C}_{N}$ using the Monte Carlo procedure described at the end of Section III.2.2 and for which the errors are generated from a standard Gaussian. Figure 2 shows that regardless of the error structure or model considered, the graphs of the simulated null distributions overlap for both the Kolmogorov-Smirnov and the Cramér-Von Mises statistics. In contrast to Figure 1, where the curves for M1 and M2 differ because the classical processes $w_{k,N}$ converge to Brownian motions with a common mean but model-dependent variance, the transformed processes $v_{k,N}$ share the same mean and variance across models, leading to overlapping null distribution of the test statistics considered. This confirms that the proposed tests are, for sufficiently large $N$, distribution-free, and the proposed Monte Carlo scheme accurately approximates the limiting null distributions of the test statistics based on $v_{k,N}$. The time required by the latter amounts to 3.74 seconds. All computations were performed on a MacBook Pro (16-inch, 2019) equipped with a 2.3 GHz 8-core Intel Core i9 processor, 16 GB of 2667 MHz DDR4 memory, and an Intel UHD Graphics 630 (1536 MB) card.
@@ III.3.2 Statistical properties
To ensure that the probability of a false rejection [^9] A false rejection means the statistical test concludes that the hypothesized model does not fit the data well even though the model is correct. – known in statistics as the probability of Type I error – does not exceed the pre-determined significance level, we generate $10^{4}$ datasets from the additive combination of the mean function $\bm{A}^{M1}_{{f_{b}},\ell}(\bm{\theta})$ and the Gaussian error $\bm{u}^{(1)}_{{f_{b}},s}$ in Eqs. (41)-(42). We test the validity of $\bm{A}^{M1}_{{f_{b}},\ell}(\bm{\theta})$ by means of the statistics $\widetilde{K}_{N}$ and $\widetilde{C}_{N}$. Table 1 shows the simulated probabilities of rejecting $\bm{A}^{M1}_{{f_{b}},\ell}(\bm{\theta})$ when choosing different significance levels. The results indicate that the simulated Type I error probabilities closely match the nominal levels.
@@
A Monte Carlo simulation involving $10^{4}$ replicates is performed for each of the models in Eq. (63) and combined with the Gaussian error structure in Eq. (42). The simulated power of the Kolmogorov–Smirnov and Cramér–von Mises test statistics are reported in Tables 2 and 3, respectively, for different significance levels $\alpha$. For both test statistics, the simulated statistical power is close to one when the significance level is 0.05 and can be as high as $0.776$ even at smaller significance levels $\alpha=0.001$.
@@ IV Testing spectral models for the SGWB
We now apply the methodology described in Sec. III to the case of a realistic astrophysical SGWB model and the data obtained by the Advanced LIGO and Advanced Virgo GW detectors in their third observing run (O3). O3 includes data from both Advanced LIGO detectors as well as the Advanced Virgo detector, collected from April 1, 2019, to March 27, 2020, with a month-long pause in October 2019. Aside from this pause, the duty cycles of the three detectors were 77%, 75%, and 76% for LIGO Livingston, LIGO Hanford, and Virgo, respectively [56]. For this study, we subdivide the O3 data set into S=S=15 segments, each corresponding to approximately 10 days of detector live time. This choice was made to ensure that each part of the sidereal day included data in each of the 15 segments. We also divide the frequency range 20-180 Hz into $B=8$ discrete frequency bins, each 20 Hz wide. We assume the same time-domain and frequency-domain data quality cuts that were applied during other O3 anisotropic SGWB search analyses [7].
@@
$A_{\ell}(f;\bm{\theta})$ $=$ $\frac{2}{\pi}\int dkk^{2}\,|\delta\Omega_{\text{GW}\,,{\ell}}(k,f;\bm{\theta})|^{2},\hbox{ and}$ (66)
$\delta\Omega_{\text{GW}\,,\ell}(k,f;\theta)$ $=$ $\frac{f}{4\pi\rho_{c}}\int dr\,\mathcal{B}(f,r;\bm{\theta})\,\big[b(r)\,\delta_{m,k}(r)j_{\ell}(kr)\big],$
@@
$\mathcal{B}(f,z;\bm{\theta})=A_{\text{\rm max}}\,f^{-1/3}\,e^{-(z-z_{c})^{2}/2\sigma_{z}^{2}}\,$, (68)
with the free parameters $\theta=(A_{\text{\rm max}},z_{c},\sigma_{z})$.
@@
$\mathcal{K}(f)=\frac{2\pi^{2}}{3H_{0}^{2}}f_{\rm ref}^{3}\left(\frac{f}{f_{\rm ref}}\right)^{\alpha}$ (72)
Here, $f_{\rm ref}$ is the reference frequency of 25Hz used in O3 [7], and $f$ again takes on values of the midpoints of each frequency band $\bar{{f_{b}}}$. We also apply the square of this scaling factor to the inverted Fisher matrix in order to obtain a properly scaled covariance matrix for the $A_{\ell}$ (Eq. (17)). For notational simplicity, for the remainder of this section, when we write ${a}_{{f_{b}},\ell m}^{M}$, we assume it to be the version scaled by $\tau_{{f_{b}},\ell}$, and when we write $\hat{a}_{{f_{b}},s,\ell m}$ or $(\Gamma^{-1}_{{f_{b}},s})_{\ell m,\ell^{\prime}m^{\prime}}$, we assume them to be the versions scaled by $\mathcal{K}(f)$ and $\mathcal{K}(f)^{2}$ respectively.
With the model map elements ${{a}}_{{f_{b}},\ell m}^{M}$ and data map elements $\hat{{a}}_{{f_{b}},s,\ell m}$ properly scaled, we may find their sum ${\hat{a}}_{{f_{b}},s,\ell m}^{\rm inj}$ to complete the signal injection. The corresponding angular power spectrum ${\hat{A}}_{{f_{b}},s,\ell}^{\rm inj}$ is calculated according to Eq. (16),
$\hat{A}^{\rm inj}_{{f_{b}},s,\ell}=\frac{1}{1+2\ell}\sum_{m=-\ell}^{\ell}\left[|\hat{a}^{\rm inj}_{{f_{b}},s,\ell m}|^{2}-(\Gamma^{-1}_{{f_{b}},s})_{\ell m,\ell m}\right],$ (73)
and the covariance matrix is computed according to Eq. (17), with ${a}_{{f_{b}},\ell m}$ replaced by $\hat{a}^{\rm inj}_{{f_{b}},\ell m}$, which is defined as the average of $\hat{a}^{\rm inj}_{{f_{b}},s,\ell m}$ over the data segments $s$.
The necessary inputs for the procedure described in Sec. III are the data ${\hat{A}}_{{f_{b}},s,\ell}^{\rm inj}$, their corresponding covariance matrices, and model values ${A}_{\ell}(\bar{{f_{b}}};\bm{\theta})$. In this study, we test our procedure on two variations of these quantities.
We compare the model described by Eqs. (66)-(68) to data injected using the same model, applying the covariance matrix as described above. When generating the injected signal, the parameter values are $1.2\times 10^{-28}$, $0.6$, and $0.7$ for $A_{max}$, $z_{c}$, and $\sigma_{z}$ respectively. The parameter estimates obtained via Eq (21) are $1.19\times 10^{-28}$, $0.590$, and $0.694$, with corresponding standard errors $5.30\times 10^{-31}$, $0.014$, and $0.012$. Next, we use the same covariance matrix and again use a postulated model as given in Eq. (68), but in this case we use data which has been injected with a different variation of the model. In particular, we replace the astrophysical kernel with
@@
where $\bm{\theta}=(\beta_{0},\beta_{1})$. For this alternative model, we set $\beta_{0}=1.2\times 10^{-28}$ and $\beta_{1}=6.0\times 10^{-29}$ when generating the injected data. The parameter estimates for the postulated model are then $2.84\times 10^{-28}$, $3.58$, and $2.72$ for $A_{max}$, $z_{c}$, and $\sigma_{z}$, respectively, with corresponding standard errors $8.29\times 10^{-30}$, $0.138$, and $0.060$.